Lilian9667:

Determine the x- and y-intercepts of the graph of x−4y=8 . Then plot the intercepts to graph the equation.

Lilian9667:

1 attachment
Lilian9667:

*I need Help* On this

Shadow:

Start by isolating y \[x - 4y = 8\] Subtract x from both sides \[-4y = -x + 8\] Divide both sides by -4 \[y = \frac{ 1 }{4 } x -2\]

Shadow:

-2 is the y intercept, you can get the x intercept by setting y to 0 then solving for x, or graphing the line.

Shadow:

Keep in mind that an intercept is when the line crosses the x or y axis.

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow Start by isolating y \[x - 4y = 8\] Subtract x from both sides \[-4y = -x + 8\] Divide both sides by -4 \[y = \frac{ 1 }{4 } x -2\] \(\color{#0cbb34}{\text{End of Quote}}\) Ok

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow Keep in mind that an intercept is when the line crosses the x or y axis. \(\color{#0cbb34}{\text{End of Quote}}\) Like a Slope?

Shadow:

The line is the slope, yes

Lilian9667:

Is this the Right Answer?

1 attachment
Shadow:

No, since -2 is the y intercept, the line crosses the y axis at -2

Shadow:

the x intercept should have some positive value since the slope is also positive and the y int is negative